A 10 g bullet moving with a velocity of 500 m/s enters a stationary piece of ice mass 10 kg and stops . If the the piece of ice is lying on a frictionless plane then it's velocity will be?


1
Expert's answer
2021-09-09T18:10:14-0400

let mass of the bullet "=m_1"

mass of ice "=m_2"

initial momentum of the bullet"=m_1v_1"

initial mass of ice =0

from conservation of momentum,

total initial momentum = total final momentum

"m_1v_1=(m_1+m_2 )v"

"\frac{10}{1000}\times 500=(\frac{10}{1000}+10)v"

"5=10.01v"

"v=0.4995m\/s" (velocity of ice)